A small stone is thrown away into the air with a velocity of 15 m/s at an angle of 30 degree along the horizontal. (With solution)


a. Find the maximum height reached by the object,

b. Time of flight, and

C. Range




1
Expert's answer
2021-11-23T10:30:57-0500

a. The height:


"h=\frac{v^2\sin^2\theta}{2g}=2.87\text{ m}."

b. Time of flight:


"t=2\frac{v\sin\theta}{g}=1.53\text{ s}."

c. Range:


"r=\frac{v^2\sin(2\theta)}{g}=19.9\text{ m}."